# Sample and Map from a Single Convex Potential: Generation using Conjugate Moment Measures

### v1

@@ 3.2 Adding a Conjugate
Let $\rho$ be an absolutely continuous probability measure supported on a compact, convex set. Then, there exists a convex function $w$ such that
$\rho=\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}.$
Similarly to the terminology we used for moment measures, we refer to $\rho$ as the conjugate moment measure of $w$, the potential $w$ as the conjugate moment potential of $\rho$, and $\mathfrak{P}_{w}$ as $\rho$’s conjugate Gibbs factor.
Proof Sketch. The complete proof, along with the theorems used in the proof, can be found in the Appendix. We rely on the identity $\nabla w^{*}\circ\nabla w=\mathrm{id}$ that holds for strictly convex functions $w$ and show the existence of a strictly convex potential $w$ that verifies $\nabla w\,\sharp\,\rho=\mathfrak{P}_{w}$. To proceed, we first introduce some notations. Let $\mathcal{L}(\mathbb{R})$ be the set of functions mapping $\mathbb{R}^{d}$ to {$\mathbb{R}\cup\{+\infty\}$. Given a non-negligible set $\Omega\subset\mathbb{R}^{d}$ and a continuous function $v:\Omega\rightarrow\mathbb{R}$ such that $\int_{\mathbb{R}^{d}}e^{-v(z)}\mathrm{d}z$ is finite, we denote by $\mathfrak{P}_{v}^{\Omega}$ the probability measure with density
$\frac{e^{-v(x)}1_{x\in\Omega}}{\int_{\Omega}e^{-v(z)}\mathrm{d}z}\,.$
We then define, for a probability $\rho$ supported on a non-negligible set $\Omega$, the map $G_{\rho}^{\Omega}:\mathcal{L}(\mathbb{R})\rightarrow\mathcal{L}(\mathbb{R})$ which to a potential $v$ associates the Brenier potential, defined in equation 4, from $\rho$ to $\mathfrak{P}_{v}^{\Omega}$
$G_{\rho}^{\Omega}(v):=\mathfrak{B}(\rho,\mathfrak{P}_{v}^{\Omega})$
Theorem 2.
$\mathcal{M}$ =$=\{f\in\mathscr{C}(\Omega)\;\text{such that}\;\;\forall x,y\in\Omega,$
$|f(x)-f(y)|\leq R\|x-y\|_{2}\;\text{and}\;f(0_{\mathbb{R}^{d}})=0\}$
3.3 Illustrations
We focus on the case where $\rho$ is Gaussian to verify that our conjugate factorization has better properties. Recall that we saw in subsection 3.1 that Gibbs factors of $\mathcal{N}(0,\Sigma)$ were multivariate normal distributions with covariance matrix $\Sigma^{-1}$.
@@ 3.3 Illustrations
When $\Sigma$ is non degenerate, $\rho=\mathcal{N}(m,\Sigma)$ is the conjugate moment measure of
$w(x)=\frac{1}{2}(x-r)^{T}\Sigma^{-1/3}(x-r),$
@@
whose Gibbs distribution is $\mathfrak{P}_{w}=\mathcal{N}(r,\Sigma^{1/3})$ where $r=(I_{d}+\Sigma^{1/3})^{-1}m$. Potential $w$ is the unique potential for which the Gibbs distribution is Gaussian, satisfying the condition $\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}=\rho\,.$
@@
The proof is similar to that of Proposition 1 and is given in the Appendix. While this potential $w$ may not be the only conjugate moment potential, the Gibbs distribution associated with it, $\mathcal{N}(r,\Sigma^{1/3})$, appears to be better suited to the target distribution $\rho$. This is particularly evident in the example shown in Figure 1. Additionally, it is worth noting that for the Gaussian case, our factorization still works when $\rho$ is not centered, whereas the appraoch of Cordero-Erausquin and Klartag (2015) requires $\rho$ to be mean 0.
@@
Figure 2: Comparison between the Gibbs factor $\mathfrak{P}_{u}$ and the conjugate Gibbs factor $\mathfrak{P}_{w}$ for two mixtures of 1D Gaussian distributions, $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$. The density plots overlay the (conjugate) Gibbs factor with $\bm{\rho}$ and a standard Gaussian distribution $\bm{\mathcal{N}(0,1)}$ for reference. Plots (a) and (c) show that the usual Gibbs factor’s spread is inverse to that of $\rho$. In plots (b) and (d) we see the far more suitable behavior of our conjugate moment factorization.
4 Estimating Conjugate Moment Measures from Data Samples
We first explain how to sample from $\rho$ when its conjugate moment decomposition is known using the Langevin Monte Carlo (LMC) algorithm and a conjugate solver. We then describe a method to estimate a conjugate moment potential of a distribution $\rho$ using i.i.d samples $(x_{1},\dots,x_{n})\sim\rho$. In this approach, the conjugate moment potential of $\rho$ is parameterized using an input convex neural network (ICNN) $w_{\theta}$ following the architecture proposed in (Vesseron and Cuturi, 2024). The conjugate moment potential $w_{\theta}$ is then estimated using an algorithm inspired by the fixed-point method associated to Theorem 2.
4.1 Sampling from a conjugate moment measure
In this paragraph, we suppose that we know a conjugate moment potential $w$ of $\rho$, i.e. $\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}=\rho$. Knowing $w$, drawing samples from $\rho$ can be done by first sampling $x_{i}\sim\mathfrak{P}_{w}$ and then applying $\nabla w^{*}$ to those points as $\nabla w^{*}(x_{i})\sim\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}=\rho$. The LMC algorithm is a widely used method for generating samples from a smooth, log-concave density like $\mathfrak{P}_{w}$ (Roberts and Tweedie, 1996; Cheng and Bartlett, 2018; Dalalyan and Karagulyan, 2019). Starting from an initial point $x^{(0)}$, the LMC algorithm iterates according to the following update rule:
$x^{(k+1)}$ $=x^{(k)}-\gamma\nabla w(x^{(k)})+\sqrt{2\gamma}z^{(k)},\;\;z^{(k)}\sim\mathcal{N}(0,I_{d}).$
where $\gamma$ is the step size. After a warm-up period, the LMC iterates are distributed according to the log-concave distribution $\mathfrak{P}_{w}$. As for the gradient of the convex conjugate $\nabla w^{*}$, it can be efficiently estimated from $w$. By applying Danskin’s envelope theorem (1966), it follows that $\nabla w^{*}(y)$ is the solution to the following concave maximization problem,
$\nabla w^{*}(y)=\arg\sup_{x}\,\langle x,y\rangle-w(x)\,.$
This optimization problem can be solved using algorithms such as gradient ascent, (L)BFGS (Liu and Nocedal, 1989), or ADAM (Kingma and Ba, 2014). Thus, having access to a conjugate moment potential $w$ of $\rho$ enables to efficiently sample from the target distribution $\rho$. When the conjugate solver is warm-started with $x_{0}=y$, gradient steps can be interpreted as a denoising procedure of the Langevin sample.
4.2 Fixed-point Approach
For a probability measure $\rho\in\mathcal{P}(\mathbb{R}^{d})$, we define the map $G_{\rho}:\mathcal{L}(\mathbb{R^{d}})\rightarrow\mathcal{L}(\mathbb{R^{d}})$ which to a potential $w$ associates the Brenier potential from $\rho$ to $\mathfrak{P}_{w}$
$G_{\rho}(w):=\mathfrak{B}(\rho,\mathfrak{P}_{w})$
The fixed point of $G_{\rho}$ correspond exactly to the conjugate moment potentials of $\rho$. This observation motivates the following fixed-point iteration scheme to compute a conjugate moment potential of $\rho$:
$w_{0}:=\tfrac{1}{2}\|\cdot\|^{2};\quad\forall t\geq 1,\quad w_{t+1}:=G_{\rho}(w_{t}).$
@@ 4.2 Fixed-point Approach
4:     Draw $y_{1},\dots,y_{n}\sim\mathfrak{P}_{w_{\theta}}$ using LMC algorithm
@@
Algorithm 1 Fixed point training of $w_{\theta}$
Starting from the standard Gaussian distribution $\mathfrak{P}_{w_{0}}=\mathcal{N}(0,I_{d})$, we iteratively compute the OT potential $w_{t+1}$ between the distribution $\rho$ and $\mathfrak{P}_{w_{t}}$ at iteration $t+1$. During the next iteration $t+2$, the updated distribution $\mathfrak{P}_{w_{t+1}}$ becomes the target distribution for computing the next OT map starting from the source $\rho$. We apply this methodology when the OT map is available in closed form at each step. In all other cases, we found that performing just one optimization step to estimate the OT map at each iteration was sufficient. For this reason, our final methodology relies on a single ICNN $w_{\theta}$ which serves two purposes at each iteration: generating samples from $\mathfrak{P}_{w_{\theta}}$ and being optimized by doing one step of the neural OT solver from Amos (2023) to approximate the OT map between $\rho$ and $\mathfrak{P}_{w_{\theta}}$. Algorithm 1 details the steps of the procedure, where $\tilde{x}(y_{i})$ is the estimation of $\nabla w_{\theta}^{*}(y_{i})$, computed as described in Amos (2023).
@@ 5 Experiments
We conduct preliminary experiments using the fixed-point iteration scheme described in Section 4.2, focusing on univariate $\rho$, for which Monge maps have a closed form. Following this, we estimate the conjugate moment potential associated with several 2D distributions as well as the MNIST digits distribution (LeCun et al., 2010).
@@ 5.1 Univariate distributions
In the 1D case, the OT map between density functions $\mu$ and $\nu$ is given by the formula (Peyré et al., 2019):
$\mathfrak{B}(\mu,\nu)=C_{\nu}^{-1}\circ C_{\mu}$
where $C_{\mu}:\mathbb{R}\rightarrow[0,1]$ is the cumulative distribution function associated to $\mu$, defined as: $C_{\mu}(x):=\int_{-\infty}^{x}\mathrm{d}\mu$. The quantile function $C_{\mu}^{-1}:[0,1]\rightarrow\mathbb{R}\cup\{-\infty\}$ is the pseudoinverse $C_{\mu}^{-1}(r):=\min\{x\in\mathbb{R}\cup\{-\infty\}:C_{\mu}(x)\geq r\}$. For a given distribution $\rho$ and an initial distribution $\mathfrak{P}_{w_{0}}$, the analytical form of the Brenier map enables the exact application of the fixed-point algorithm described in section 4.2. Further implementation details can be found in the Appendix. We use this methodology to compute the moment potential and the conjugate moment potential of two distinct mixtures of Gaussians. Display (a) in Figure 2 shows that the Gibbs factor of $\rho_{1}$ has a longer tail compared to the centered standard Gaussian. This aligns with theoretical results on Gaussians, which suggest that more concentrated distributions yield broader Gibbs factors. Conversely, for the second measure $\rho_{2}$ in (c), which has a larger variance, the Gibbs factor is more concentrated. By design, our conjugate Gibbs factor in (b) and (d) closely matches the shapes of $\rho_{1}$ and $\rho_{2}$, making it a more suitable prior for generative modeling applications. See also Figure 5 in the appendix, where we display $\nabla u\,\sharp\,\mathfrak{P}_{u}$ (for the moment potential $u$) and $\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}$ (for the conjugate moment potential $w^{*}$) with their respective measures $\rho_{1}$ and $\rho_{2}$.
@@
Measure $\rho$, potential $w_{\theta}$ $\nabla w_{\theta}^{*}\,\sharp\,e^{-w_{\theta}}$efer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a)(b)(c)(d)(e)
@@
Figure 3: Measure $\rho$ and level sets of its convex conjugate potential $w_{\theta}$ (top), and $\nabla w_{\theta}^{*}\,\sharp\,e^{-w_{\theta}}$ (bottom).
5.2 Generative Modeling
We focus on several 2D distributions as well as the MNIST dataset (LeCun et al., 2010). Our first step is to estimate a conjugate moment potential $w$ for these distributions using the fixed-point method outlined in Section 4.2. Following this, we generate new images based on the conjugate potential $w$ by utilizing the methodology detailed in Section 4.1. The hyperparameters used for these experiments are provided in the Appendix.
@@ 5.2 Generative Modeling
Figure 4: Learning of the conjugate moment measure. The single-step generative model $\nabla w_{\theta}$ converts highly structured data $\rho$ into a log-concave distribution $\mathfrak{P}_{w_{\theta}}$. The train distribution $\rho$ is sampled via optimization, in an iterative manner. Strict convexity of $w_{\theta}$ guarantees both (i) invertibility of $\nabla w_{\theta}$, and (ii) that Langevin dynamics recovers correctly the conjugate moment $\mathfrak{P}_{w_{\theta}}$.
2D Experiments. For the 2D distributions: (a) Circles, (b) S-curve, (c) Checkerboard, (d) Scaled-Rotated S-curve and (e) Diag-Checkerboard, the fixed-point approach proved effective. The results, shown in Figure 3, demonstrate that we successfully estimated the conjugate potential associated with these distributions, as the conjugate moment measure of the estimated potential is well aligned with the target distributions $\rho$. Additionally, in the 2D experiments (c), (d) and (e), we can clearly observe how the Gibbs factor closely follows the shape of the distribution $\rho$.
MNIST Experiment. The fixed-point method also performs well on the MNIST dataset. As shown in Figure 4, we successfully generated digits using our approach. To the best of our knowledge, this is the first instance where the MNIST distribution has been generated using an ICNN. The structure of the Gibbs distribution, compared to Gaussian noise, seems to facilitate the generative task of transporting $\mathfrak{P}_{w}$ to the MNIST distribution.

### v2

@@ 3.2 The Conjugate Moment Measure factorization
Let $\rho\in\mathcal{P}(\mathbb{R}^{d})$ be an absolutely continuous probability measure supported on a compact, convex set. Then, there exists a convex function $w$ such that $\rho=\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}.$
Accordingly, we now refer to $\rho$ as the conjugate moment measure of $w$, $w$ as the conjugate moment potential of $\rho$, and $\mathfrak{P}_{w}$ as the conjugate Gibbs factor of $\rho$. Our proof strategy differs significantly from that used in (Cordero-Erausquin and Klartag, 2015) and (Santambrogio, 2016) since we use Schauder’s fixed point theorem to show that the following map admits a fixed point:
$G_{\rho}:\mathcal{L}(\mathbb{R}^{d})\rightarrow\mathcal{L}(\mathbb{R}^{d})\,,\quad G_{\rho}(w):=\mathfrak{B}(\rho,\mathfrak{P}_{w})$ (7)
where $\mathcal{L}(\mathbb{R}^{d})$ is the set of functions mapping $\mathbb{R}^{d}$ to {$\mathbb{R}\cup\{+\infty\}$, and $G_{\rho}$ assigns to a potential $w$ the Brenier potential transporting $\rho$ to $\mathfrak{P}_{w}$. Fixed points of $G_{\rho}$ correspond precisely to the conjugate moment potentials of $\rho$, and their existence guarantees a solution to the moment measure factorization. A sketch of the proof is provided in Appendix A.1, with full details in Appendix A.2. When $\rho$ is Gaussian, the conjugate moment measure factorization admits an explicit solution, as given in Proposition 2. Recall that we saw in §3.1 that Gibbs factors of $\mathcal{N}(0,\Sigma)$ were multivariate normal distributions with covariance matrix $\Sigma^{-1}$.
@@
When $\Sigma$ is non-degenerate, $\rho=\mathcal{N}(m,\Sigma)$ is the conjugate moment measure of $w(x)=\frac{1}{2}(x-r)^{T}\Sigma^{-1/3}(x-r),$ whose Gibbs distribution is $\mathfrak{P}_{w}=\mathcal{N}(r,\Sigma^{1/3})$ where $r=(I_{d}+\Sigma^{1/3})^{-1}m$. The function $w$ is the unique conjugate moment potential of $\rho$ whose Gibbs distribution remains Gaussian.
@@
The proof is provided in Appendix A.4 and relies on the fact that the OT map between two Gaussians is known in closed form Peyré et al. (2019). While this potential $w$ may not be the only conjugate moment potential, the Gibbs distribution associated with it, $\mathcal{N}(r,\Sigma^{1/3})$, appears to be better suited to the target distribution $\rho$. This is particularly evident in the example shown in Figure 1. Additionally, it is worth noting that for the Gaussian case, our factorization still works when $\rho$ is not centered, whereas the approach of Cordero-Erausquin and Klartag (2015) requires $\rho$ to be mean 0.
3.3 Monge-Ampère Equation for Conjugate Moment Measures
From the moment measure factorization, $\rho=\nabla u\sharp\mathfrak{P}_{u},$ which holds at the level of probability distributions, one can derive the corresponding equality between probability density functions and obtain the following Monge–Ampère equation (Cordero-Erausquin and Klartag, 2015):
$\rho(x)=e^{-\mathcal{E}_{u}(x)}\quad\text{with}\quad\mathcal{E}_{u}(x)=u(\nabla u^{*}(x))-\ln(\det H_{u^{*}}(x))+\ln(C_{u}),$ (8)
where $u^{*}$ is the convex conjugate of $u$, $H_{u^{*}}(x)$ denotes the Hessian of $u^{*}$ at $x$, and $C_{u}=\int e^{-u}$ is the normalizing constant ensuring that $\mathfrak{P}_{u}=e^{-u}/C_{u}$ is a probability distribution. Similarly, Theorem 1 guarantees the existence of a convex function $w$ that satisfies the following Monge-Ampère equation for any absolutely continuous probability measure $\rho$ supported on a compact, convex set:
$\rho(x)=e^{-\mathcal{E}_{w}(x)}\quad\text{where}\quad\mathcal{E}_{w}(x)=w(\nabla w(x))-\ln(\det H_{w}(x))+\ln(C_{w}),$ (9)
with $H_{w}(x)$ denoting the Hessian of $w$ at $x$ and $C=\int e^{-w}$ being the normalizing constant of $\mathfrak{P}_{w}$. The potentials $w$ that solve (9) are precisely the conjugate moment potentials of $\rho$.
4 Estimating Conjugate Moment Potentials in Practice
We first explain how to sample from $\rho$ when one of its conjugate moment potential is known, using the Langevin Monte Carlo (LMC) algorithm and a conjugate solver. We then describe a method to estimate a conjugate moment factorization of $\rho$ using i.i.d samples $(x_{1},\dots,x_{n})\sim\rho$. In this approach, the conjugate moment potential of $\rho$ is parameterized using an input convex neural network (ICNN) $w_{\theta}$ following the architecture proposed in (Vesseron and Cuturi, 2024). The conjugate moment potential $w_{\theta}$ is then estimated using an algorithm inspired by the fixed-point method associated to Theorem 1. Finally, we address the case where the density of $\rho$ is known up to a normalizing constant, a common scenario in sampling; we use an ICNN to parameterize the potential and estimate it via regression using the Monge–Ampère equation (9).
4.1 Sampling from $\rho$ using its Conjugate Moment Factorization
In this paragraph, we suppose that we know a conjugate moment potential $w$ of $\rho$, i.e. $\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}=\rho$. Knowing $w$, drawing samples from $\rho$ can be done by first sampling $x\sim\mathfrak{P}_{w}$ and then applying $\nabla w^{*}$ to those points as $\nabla w^{*}(x)\sim\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}=\rho$. The LMC algorithm is a widely used method for generating samples from a smooth, log-concave density like $\mathfrak{P}_{w}$ (Roberts and Tweedie, 1996; Cheng and Bartlett, 2018; Dalalyan and Karagulyan, 2019). Starting from an initial point $x^{(0)}$, the LMC algorithm iterates according to the following update rule:
$x^{(k+1)}=x^{(k)}-\tau\nabla w(x^{(k)})+\sqrt{2\tau}z^{(k)},\;\;z^{(k)}\sim\mathcal{N}(0,I_{d}),$
where $\tau$ is the step size. After a warm-up period, the LMC iterates are distributed according to the log-concave distribution $\mathfrak{P}_{w}$. As for the gradient of the convex conjugate $\nabla w^{*}$, it can be efficiently estimated from $w$. By applying Danskin’s envelope theorem (1966), it follows that $\nabla w^{*}(y)$ is the solution to the following concave maximization problem:$\nabla w^{*}(y)=\arg\sup_{x}\,\langle x,y\rangle-w(x)\,.$ This optimization problem can be solved using algorithms such as gradient ascent, (L)BFGS (Liu and Nocedal, 1989), or ADAM (Kingma and Ba, 2014). Thus, having access to a conjugate moment potential $w$ of $\rho$ enables to efficiently draw samples from it. Note that the gradient steps of the conjugate solver can be interpreted as a denoising procedure applied to samples drawn from $\mathfrak{P}_{w}$.
4.2 Learning the Conjugate Moment Factorization from Samples: CMFGen
We consider the case where $\rho\in\mathcal{P}(\mathbb{R}^{d})$ can only be accessed through samples, as in an empirical distribution approximation $\rho_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$. The fixed points of the map $G_{\rho}$, as defined in (7), correspond exactly to the conjugate moment potentials of $\rho$. This observation motivates the following fixed-point iteration scheme to compute a conjugate moment potential of $\rho$:
$w_{0}:=\tfrac{1}{2}\|\cdot\|^{2};\quad\forall t\geq 1,\quad w_{t+1}:=G_{\rho}(w_{t}),$ (10)
Starting from $w_{0}=\tfrac{1}{2}\|\cdot\|^{2}$, which corresponds to $\mathfrak{P}_{w_{0}}=\mathcal{N}(0,I_{d})$, we iteratively compute the Brenier potential $w_{t+1}$ between the distribution $\rho$ and $\mathfrak{P}_{w_{t}}$ at iteration $t+1$. During the next iteration $t+2$, the updated distribution $\mathfrak{P}_{w_{t+1}}$ becomes the target distribution to compute the next Brenier potential starting from the source $\rho$. This process is repeated until the algorithm converges.
@@ 4.2 Learning the Conjugate Moment Factorization from Samples: CMFGen
Figure 2: Comparison between the Gibbs factor $\mathfrak{P}_{u}$ and the conjugate Gibbs factor $\mathfrak{P}_{w}$ for two mixtures of 1D Gaussian distributions, $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$. The density plots overlay the (conjugate) Gibbs factor with $\bm{\rho}$ and a standard Gaussian $\bm{\mathcal{N}(0,1)}$ for reference. Gibbs factors spread inversely to $\rho$ ((a), (c)) while conjugate Gibbs factors show more suitable alignment ((b), (d)).
Univariate distributions
In the 1D case, the OT map between probability density functions $\mu$ and $\nu$ has a closed-form expression given by (Peyré et al., 2019):
$\nabla\mathfrak{B}(\mu,\nu)=C_{\nu}^{-1}\circ C_{\mu}$
where $C_{\mu}:\mathbb{R}\rightarrow[0,1]$ is the cumulative distribution function associated to $\mu$, defined as: $C_{\mu}(x):=\int_{-\infty}^{x}\mathrm{d}\mu$. The quantile function $C_{\mu}^{-1}:[0,1]\rightarrow\mathbb{R}\cup\{-\infty\}$ is the pseudoinverse $C_{\mu}^{-1}(r):=\min\{x\in\mathbb{R}\cup\{-\infty\}:C_{\mu}(x)\geq r\}$. For a given distribution $\rho$ and an initial distribution $\mathfrak{P}_{w_{0}}$, the analytical form of the Brenier map enables the execution of the algorithm in (10). To compare the conjugate moment measure factorization with the moment measure factorization in 1D, we also propose a similar 1D algorithm for estimating a moment measure potential of a distribution $\rho$, described in Appendix B.1, as no method currently exists for solving the moment measure factorization. The code for both algorithms is also provided in Appendix B.1.
Higher dimensional case
In higher dimensions, where the OT map is not available in closed form, we estimate it using ICNNs with the architecture used in (Vesseron and Cuturi, 2024) and the neural OT solver proposed in (Amos, 2023). We found that performing a single optimization step to estimate the OT map at each iteration was sufficient. As a result, our final methodology relies on a single ICNN, $w_{\theta}$, which serves two purposes at each iteration: generating samples from $\mathfrak{P}_{w_{\theta}}$ and being optimized by a single step of the neural OT solver to approximate the OT map between $\rho$ and $\mathfrak{P}_{w_{\theta}}$. Algorithm 1 details the steps of the procedure, where $\tilde{x}(y_{i})$ is the estimation of $\nabla w_{\theta}^{*}(y_{i})$, computed as described in (Amos, 2023). The LMC sampling step in step 4 of Algorithm 1 requires selecting the step size hyperparameter $\tau$, which depends on $w_{\theta}$.
Algorithm 1 CMFGen algorithm
@@
4:     Draw $y_{1},\dots,y_{n}\sim\mathfrak{P}_{w_{\theta}}$ using LMC
@@
Since $w_{\theta}$ evolves during the algorithm, the step size must be dynamically adjusted. Following Proposition 1 in Dalalyan (2016), we set $\tau=\frac{1}{M_{\theta}}$, where $M_{\theta}$ is the largest eigenvalue of the Hessian of $w_{\theta}$, estimated from the current minibatch $\{x_{1},\dots,x_{n}\}$. To accelerate convergence, the LMC algorithm is initialized using particles from the previous iteration.
4.3 Learning the Conjugate Moment Factorization from an Energy: CMFMA
We now consider the case where the density of $\rho$ is known only up to a normalizing constant, which is the typical setting in sampling-based frameworks. Specifically, we assume access to an energy function $\mathcal{E}$ such that $\rho\propto e^{-\mathcal{E}}$. To learn a conjugate moment potential of $\rho$, we propose to leverage the Monge–Ampère formulation defined in (9), and parameterize the potential using an ICNN $w_{\theta}$ trained via regression on the following objective:
$\mathbb{E}_{x\sim\mathbb{P}}\left[\|\mathcal{E}(x)-w_{\theta}(\nabla w_{\theta}(x))+\ln(\det H_{w_{\theta}}(x))\|^{2}\right].$
In this objective, $\mathbb{P}$ can be taken as the uniform distribution over the sampling space when the dimension is small and the domain is bounded. In higher-dimensional settings, $\mathbb{P}$ can instead be chosen as $\rho$, with samples obtained via the Langevin Monte Carlo (LMC) algorithm. At the end of training, the learned potential $w_{\theta}$ satisfies $w_{\theta}(\nabla w_{\theta}(x))+\ln(\det H_{w_{\theta}}(x))\approx\mathcal{E}(x)$ and the pushforward $\nabla w_{\theta}^{*}\sharp\mathfrak{P}_{w_{\theta}}$ provides a close approximation of the target $\rho$. §4.1 details how to sample from $\nabla w_{\theta}^{*}\sharp\mathfrak{P}_{w_{\theta}}$ using only the learned potential $w_{\theta}$. Note that, in contrast, learning the moment potential $u$ via regression seems particularly challenging due to the presence of the Legendre transform $u^{*}$ in multiple terms of the Monge–Ampère equation (8).
@@ 5 Experiments
We begin with preliminary experiments using the CMFGen and CMFMA algorithms introduced in Section 4.2 and 4.3, starting with univariate distributions $\rho$ for which the Monge map is available in closed form. We then estimate the conjugate moment potential for several 2D distributions, either from samples (using CMFGen) or from an energy function (using CMFMA). Finally, we demonstrate the scalability of CMFGen by applying it to higher-dimensional datasets such as MNIST (LeCun et al., 2010) and the Cartoon dataset from Royer et al. (2018).
@@ 5.1 Univariate distributions
We compute the conjugate moment potential for two univariate Gaussian mixtures, $\rho_{1}$ and $\rho_{2}$, shown in Figure 2, using both CMFGen and CMFMA. CMFGen uses i.i.d. samples from the mixture, while CMFMA leverages the log-density. The distribution $\rho_{1}$ is sharply concentrated around zero, whereas $\rho_{2}$ exhibits heavier tails. For comparison, we also compute the standard moment potentials using the fixed-point method described in Appendix B.1. The pushforward densities $\nabla w^{*}\sharp\mathfrak{P}_{w}$ and $\nabla u\sharp\mathfrak{P}_{u}$ (Figures 11 and 8) closely match the target distributions, confirming that the three algorithms successfully recover the (conjugate) moment potentials for $\rho_{1}$ and $\rho_{2}$. The recovered conjugate potentials are shown in Figures 9 and 10. As illustrated in Figure 2 (a), the Gibbs factor associated with $\rho_{1}$ has heavier tails than the standard Gaussian, consistent with the theoretical insight that concentrated distributions yield broader Gibbs measures. Conversely, the broader $\rho_{2}$ induces a more concentrated Gibbs factor (panel (c)). In contrast, the conjugate Gibbs factors (panels (b) and (d)) more closely match the target distributions.
Figure 3: Learning the conjugate moment potential from an energy. $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ are learned by regression with CMFMA. The second column shows the learned energy; the third displays the corresponding conjugate moment potential; the fourth shows samples (in red) drawn from $\nabla w_{\theta}^{*}\sharp\mathfrak{P}_{w_{\theta}}$.
5.2 2D Experiments
We consider several 2D distributions defined either through samples—(a) Circles, (b) S-curve, (c) Checkerboard, (d) Scaled-Rotated S-curve, and (e) Diag-Checkerboard (Figure 4)—or through known energy functions $\mathcal{E}_{1}$, $\mathcal{E}_{2}$, and $\mathcal{E}_{3}$ (Figures 3and 12). Our first step is to estimate a conjugate moment potential $w_{\theta}$ for these distributions using either the CMFGen or the CMFMA algorithm. Following this, we generate new samples based on the learned conjugate potential $w_{\theta}$ using the methodology detailed in §4.1. The potential $w_{\theta}$ is implemented as an ICNN with five hidden layers of size $128$ and quadratic input connections, based on the architecture of Vesseron and Cuturi (2024). Detailed hyperparameters for both 2D and high-dimensional experiments are provided in Appendix E.
CMFMA.   The energy functions $\mathcal{E}_{1}$, $\mathcal{E}_{2}$, and $\mathcal{E}_{3}$ are standard 2D benchmarks for evaluating optimization algorithms; their analytical forms are provided in Appendix C.2. For all three, the ICNNs $w_{\theta}$ successfully learn the corresponding energy landscape and permit to draw new samples from the target distributions, as demonstrated in Figure 3 and 12.
@@
Measure $\rho$ potential $w_{\theta}$$\nabla w_{\theta}^{*}\,\sharp\,\mathfrak{P}_{w_{\theta}}$efer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a)(b)(c)(d)(e)
@@
Figure 4: Samples from $\rho$ (top), level sets of $w_{\theta}$ (middle) and samples from $\nabla w_{\theta}^{*}\,\sharp\,e^{-w_{\theta}}$ (bottom).
CMFGen.   Our method accurately estimates the conjugate moment potentials, as the associated measures $\nabla w_{\theta}^{*}\,\sharp\,\mathfrak{P}_{w_{\theta}}$ closely align with the target distributions in Figure 4. The second row further illustrates that the conjugate Gibbs factor follows the shape of the distribution $\rho$ in each case. For comparison, we train an ICNN to map a Gaussian directly to the target distribution using the solver of Amos (2023). The boxplots in Figure 13, which show the Sinkhorn divergence (Ramdas et al., 2017) between generated and target data, demonstrate that CMFGen consistently produces samples of equal or superior quality. Note that CMFGen introduces no additional hyperparameters compared to the generative ICNN, aside from the number of LMC steps used when sampling from $\mathfrak{P}_{w_{\theta}}$.
5.3 High-dimensional experiments.
Image generation.
Refer to caption     Refer to caption     Refer to caption
Figure 5: MNIST Generation using CMFGen. Samples from the Gibbs noise distribution $\mathfrak{P}_{w_{\theta}}$ (left); digits generated from $\nabla w_{\theta}^{*}\sharp\mathfrak{P}_{w_{\theta}}$ (middle); digits generated by an ICNN trained to directly transport Gaussian noise to MNIST (right).
We evaluate CMFGen on MNIST (LeCun et al., 2010) and the Cartoon dataset (Royer et al., 2018). As illustrated in Figures 5, and 6, (see also 14 for additional generated cartoons), CMFGen successfully generates visually coherent images of digits and cartoon faces. Interestingly, the noise sampled from the log-concave distribution $\mathfrak{P}_{w_{\theta}}$ already exhibits features of the generated distribution: digits and faces emerge directly in the noise. To the best of our knowledge, this is the first instance where the MNIST distribution has been successfully generated using an ICNN. For comparison, Figures 5 and 15 show samples generated by ICNNs trained to transport a Gaussian distribution to the MNIST and Cartoon data. For both datasets, CMFGen generates samples of higher quality.
@@
Figure 6: Cartoon Generation using CMFGen. The conjugate potential $w_{\theta}$ is parameterized as an ICNN following Vesseron and Cuturi (2024), with five hidden layers of size 512 and four quadratic input connections, each with two additional layers of size 512. After training with CMFGen, the generative map $\nabla w_{\theta}$ transforms structured data $\rho$ into the log-concave distribution $\mathfrak{P}_{w_{\theta}}$. To sample from $\rho$, we first draw a sample from $\mathfrak{P}_{w_{\theta}}$ using the LMC algorithm, and then apply a conjugate solver to iteratively invert the map $\nabla w_{\theta}$. The strict convexity of $w_{\theta}$ ensures both (i) invertibility of $\nabla w_{\theta}$, and (ii) correctness of Langevin dynamics for sampling from $\mathfrak{P}_{w_{\theta}}$.
Image reconstruction.
Figure 7: Image inpainting on Cartoon.
Similar to normalizing flows (Rezende and Mohamed, 2015), we have access to the (unnormalized) probability density of the distribution $\nabla w_{\theta}^{*}\sharp\mathfrak{P}_{w_{\theta}}$ generated by CMFGen and CMFMA: that density is proportional to $e^{-\mathcal{E}_{w_{\theta}}(x)}$ where $\quad\mathcal{E}_{w_{\theta}}(x)={w_{\theta}}(\nabla{w_{\theta}}(x))-\ln(\det H_{{w_{\theta}}}(x))$ (see §3.3). After training $w_{\theta}$, this enables downstream tasks such as image in-painting. To evaluate this, we mask half of the pixels in test samples from MNIST and the Cartoon dataset, and perform gradient ascent on masked pixels to maximize the log-probability of the full image. As seen in Figures 7,  16 and 17, our method effectively reconstructs missing regions.